# LightPAL: Lightweight Passage Retrieval for Open Domain Multi-Document Summarization

### v1

@@ Appendix A Prompts for Generation/Evalaution
A.1 Prompts for pair-wise summary comparison
You will be given a question and two answers (A/B) written for the question.
Your task is to select answer A or B based on their relevance.
Please make sure you read and understand these instructions carefully. Keep this document open while reviewing, and refer to it as needed.
Your response must be just "A" or "B".
@@ A.1 Prompts for pair-wise summary comparison
Relevance - Selection of important content from the source. The answer should include only important information regarding the question from the source text. Annotators were instructed to penalize answers which contained redundancies and excess information.
@@
Answer A:
{{AnswerA}}
Answer B:
{{AnswerB}}
Evaluation Form ("A"/"B" ONLY):
Please indicate your choice by entering "A" or "B".
Your response:
@@
You will be given a question and two answers (A/B) written for the question.
Your task is to select answer A or B based on their diversity.
Please make sure you read and understand these instructions carefully. Keep this document open while reviewing, and refer to it as needed.
Your response must be just "A" or "B".
@@
Diversity - How varied and insightful is the answer in providing different perspectives on the question? Assess whether the answer introduces multiple viewpoints and facilitates a broader understanding of the topic.
@@
Answer A:
{{AnswerA}}
Answer B:
{{AnswerB}}
Evaluation Form ("A"/"B" ONLY):
Please indicate your choice by entering "A" or "B".
Your response:
@@
You will be given a question and two answers (A/B) written for the question.
Your task is to select answer A or B based on their comprehensiveness.
Please make sure you read and understand these instructions carefully. Keep this document open while reviewing, and refer to it as needed.
Your response must be just "A" or "B".
@@
Comprehensiveness - How thorough and detailed is the answer in covering all aspects of the question?
@@
Answer A:
{{AnswerA}}
Answer B:
{{AnswerB}}
Evaluation Form ("A"/"B" ONLY):
Please indicate your choice by entering "A" or "B".
Your response:
@@ A.2 Prompts for summary generation
You are a helpful assistant that gives long answer to question based on long stories. Write an answer based on the following question and the separated sections in the stories. Please write the answer in approximately {{n_summ_words}} words.
@@
You are a helpful assistant that gives long answer to question based on long meetings. Write an answer based on the following question and the separated sections in the meetings. Please write the answer in approximately {{n_summ_words}} words.
@@
You are a helpful assistant that gives long answer to question based on documents. Write an answer based on the following question and the documents. Please write the answer in approximately {{n_summ_words}} words.
@@
Listing 6: Prompt for summary generation in QsumOD dataset
Appendix B Implementation Details
B.1 Used packages
We used rustworkx  [^9] https://github.com/Qiskit/rustworkx version 0.14.2 for evaluating Personalized PageRank on a CPU machine. For evaluation on a GPU machine, we used cuGraph  [^10] https://github.com/rapidsai/cugraph version 24.4.0. We used vllm  [^11] https://github.com/vllm-project/vllm version 0.4.2 for inference to generate summaries.

### v2

@@ Appendix A Prompts for Generation/Evalaution
A.1 Prompts for summary evaluation
You will be given separated chunks of text and a question. You will then be given one answer written for the question based on the text.
Your task is to rate the answer on one metric.
Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed.
@@ A.1 Prompts for summary evaluation
Relevance (1-5) - selection of important content from the source. The answer should include only important information regarding the question from the source text. Annotators were instructed to penalize answers which contained redundancies and excess information.
Evaluation Steps:
1. Read the answer and the source text carefully.
2. Compare the answer to the source document and identify the main points of the source text regarding the question.
3. Assess how well the answer covers the main points of the source text regarding the question, and how much irrelevant or redundant information it contains.
4. Assign a relevance score from 1 to 5.
@@
Answer:
{{Summary}}
Evaluation Form (scores ONLY):
- Relevance:
@@
You will be given a question and one answer written for the question.
Your task is to rate the answer on one metric.
Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed.
@@
Diversity (1-5) - How varied and insightful is the answer in providing different perspectives on the question? Assess whether the answer introduces multiple viewpoints and facilitates a broader understanding of the topic.
Evaluation Steps:
1. Read and Understand: Begin by reading the question and the answer to understand the overall approach.
2. Identify Perspectives: Determine whether the answer includes multiple viewpoints, reflecting different interpretations or aspects of the topic.
3. Evaluate Depth of Insights: Assess the richness of insights provided by these viewpoints, considering how they contribute to a deeper understanding of the topic.
4. Assign a Score: Based on your evaluation, assign a diversity score from 1 to 5, reflecting the variety and depth of perspectives and insights presented in the answer.
@@
Answer:
{{Summary}}
Evaluation Form (scores ONLY):
- Diversity:
@@
You will be given a question and one answer written for the question.
Your task is to rate the answer on one metric.
Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed.
@@
Comprehensiveness (1-5) - Does the answer cover all aspect of the question? Evaluate whether the answer provides a thorough representation of information that directly addresses the question.
Evaluation Steps:
1. Read and Understand: Start by reading the question and the provided answer thoroughly to grasp the context and scope.
2. Verify Coverage: Check if the answer directly addresses all aspect of the question.
3. Assess Detail: Evaluate the detail in the explanations provided, ensuring that they are sufficient to fully address the question.
4. Assign a Score: Based on your evaluation, assign a comprehensiveness score from 1 to 5, reflecting how completely the answer addresses the question.
@@
Answer:
{{Summary}}
Evaluation Form (scores ONLY):
- Comprehensiveness:
@@ A.2 Prompts for summary generation
You are a helpful assistant that gives long answer to question based on long stories. Write an answer based on the following question and the separated sections in the stories. Please write the answer in approximately {{n_summ_words}} words. When summarizing, use only information explicitly stated in the provided story. Do not use your own knowledge or make inferences beyond what is directly stated in the text.
@@
You are a helpful assistant that gives long answer to question based on long meetings. Write an answer based on the following question and the separated sections in the meetings. Please write the answer in approximately {{n_summ_words}} words. When summarizing, use only information explicitly stated in the provided meetings. Do not use your own knowledge or make inferences beyond what is directly stated in the text.
@@
You are a helpful assistant that gives long answer to question based on documents. Write an answer based on the following question and the documents. Please write the answer in approximately {{n_summ_words}} words. When summarizing, use only information explicitly stated in the provided documents. Do not use your own knowledge or make inferences beyond what is directly stated in the text.
@@
Listing 6: Prompt for summary generation in RAG-QA dataset
Appendix B Human Annotation Procedure
Figure 4 and Figure 5 show the instructions for annotation and the annotation interface, respectively. The human annotation was conducted using exactly the same aspects (relevance, comprehensiveness, diversity) as the GPT-4 evaluation, and the descriptions for each aspect were identical to those used in the prompts for GPT-4.
Figure 4: Annotation instruction
Figure 5: Annotation interface
Appendix C Examples of Generated Summaries
Figures 6 and 7 show examples of generated summaries when LightPAL is applied to the base retriever (BM25/bge-large-en-v1.5).
Dataset: RAG-QA Query: what baseball rules or calls have become customary despite them not being in the rule book? Reference Summary: In baseball, some of the unwritten rules include not stealing bases when there is a large lead, not placing your bat on the plate following a walk, and avoiding crossing over the pitching mound on your way back to the dugout after being called out. Other unspoken rules include that one should not walk a right-handed hitter to pitch to a left-handed hitter when a right-handed pitcher is on the mound. It is also considered taboo to talk about a no-hitter while it’s in progress. Additionally, batters are encouraged to hit the ball where it is pitched. In a specific play referred to as "The neighborhood play" or sometimes "The Phantom double" it’s accepted when a fielder does not actually make contact with the base during a double play attempt. Contradicting official MLB laws is the unwritten rule that a tie goes to the runner, even though the official rulebook indicates scenarios where either the ball must arrive before the batter or vice versa. Lastly, it’s frowned upon to bunt when already holding a significant lead, as this is seen as an unnecessary strategy. BM25: Scores: pre@k: 0.4000, rec@k: 0.3333, Rel: 2.3868, Comp: 3.3359, Div: 3.0131, ROUGE: 20.2281 ✗: The concept of the "tie goes to the runner" in baseball is an example of an unwritten rule that contradicts official MLB rules. According to Rule 6.05j, a batter is out if they are tagged before reaching first base after a third strike or hitting a fair ball. However, Rule 7.08e states that any runner is out if they fail to reach the next base before being tagged. This results in a discrepancy where the outcome depends on whether the ball or the runner reaches the base first. Additionally, other unwritten rules include not walking a right-handed batter to pitch to a left-handed batter, not mentioning a no-hitter while it’s in progress, maintaining a manager’s detachment from players, and retaliating if a player is hit by a pitch. These rules, while not officially written, are widely accepted and practiced within the baseball community. BM25 + LightPAL: Scores: pre@k: 0.8000, rec@k: 0.6667, Rel: 2.8995, Comp: 4.0558, Div: 3.9707, ROUGE: 26.4306 ✓: The concept of the "tie goes to the runner" is an example of a baseball rule that is often contradictory and not explicitly stated in the official MLB rules. According to Rule 6.05j, a batter is out if they are tagged before reaching first base after a third strike or hitting a fair ball. However, Rule 7.08e states that any runner is out if they fail to reach the next base before being tagged. This creates a situation where the outcome depends on whether the ball or the runner reaches the base first, effectively giving the tie to whichever team is closer to achieving their objective first. Additionally, there are several unwritten rules in baseball that have become customary despite not being officially written in the rulebook. These include not stealing when there is a large lead, not placing the bat on the plate after a walk, and not running over the pitching mound after an out. Other examples include the "neighborhood play" at second base, where the fielder doesn’t touch the base during a double play attempt, and batters not making an effort to get out of the way of a pitch. An outfielder using the wall for support while attempting to catch a home run is also considered against the rules. These unwritten rules often revolve around sportsmanship and etiquette within the game.
Figure 6: Examples of generated summaries in RAG-QA data using BM25 and BM25+LightPAL
Dataset: Story Query: What is the dynamic between Captain Midas and Mister Spinelli in the story of CAPTAIN MIDAS? Reference Summary: Mister Spinelli is Third Officer under the command of Captain Midas and was the first to report the derelict ship and observe its potential to be claimed by the Maid. Spinelli is the first and only crew member to identify the metal from the abandoned ship as gold when he saw Captain Midas with it. The tension between Midas and Spinelli escalates and their relationship becomes antagonistic as both of them desire to benefit the most from this valuable gold and with Midas constantly pulling his authority over Spinelli. After Midas barrs him from being a part of the investigative crew, suspicion arises between the two as Spinelli suspects Midas wishes to keep the pot of gold for himself and Midas thinks that Spinelli may be telling others. This tension further escalates as Midas sees Spinelli nearly hit the trigger of the gun and in rage, the two end up fighting each other before Midas aimed a kick at his temple and killed him. … However, this relationship later sours once Spinelli finds out that the discovered ship is made out of gold. The captain is wary because Spinelli is big and dangerous, preferring to keep an eye on him in case anything goes wrong. Spinelli, on the other hand, is suspicious of the captain and goes off to tell Zaleski even though the captain said to not mention the gold to any of the rest of the crew. The captain threatens to clap Spinelli to irons, but this threat does not work because the Third Officer chooses to disobey his orders anyways. Later, Spinelli and the captain get into conflict again, with Spinelli accusing the captain that he was planning to keep all of the treasure for himself. Spinelli assumes that everybody is faking it to keep the ship, but the captain knocks him away. The two of them fight brutally, and Captain Midas has to end up killing Spinelli. bge-large-en-v1.5: Scores: pre@k: 0.4000, rec@k: 0.5000, Rel: 2.7597, Comp: 3.5871, Div: 2.1482, ROUGE: 21.6848 ✗: In the story of CAPTAIN MIDAS, the dynamic between Captain Midas and Mister Spinelli is one of suspicion and distrust. Captain Midas becomes aware of Spinelli’s intense interest in the gold found on the derelict spaceship and suspects that Spinelli may be planning to steal it for himself. To prevent this, Captain Midas decides to keep the gold a secret from the rest of the crew and assigns Spinelli to remain on the ship with him, rather than allowing him to join the prize crew that will be boarding the derelict. However, Captain Midas also notes that Spinelli seems to be watching Zaleski, suggesting that Spinelli may be trying to prevent Zaleski from double-crossing him. Despite this, Captain Midas remains wary of Spinelli and keeps a close eye on him, ultimately deciding to relieve him of his duties and assign him to stay on the ship with him. bge-large-en-v1.5 + LightPAL: Scores: pre@k: 0.8000, rec@k: 1.0000, Rel: 3.0666, Comp: 4.5831, Div: 2.8078, ROUGE: 20.0084 ✓: The dynamic between Captain Midas and Mister Spinelli in the story of CAPTAIN MIDAS is one of tension and mistrust. Initially, Captain Midas assigns Spinelli to stay on board the Maid with him, while Cohn and Zaleski take charge of the prize crew to bring the derelict ship into Callisto for salvage. However, Spinelli becomes suspicious that Captain Midas is planning to cheat him out of his rightful share of the treasure, and he begins to conspire with Zaleski to keep the treasure ship for themselves. As the story progresses, Spinelli’s actions become increasingly erratic and aggressive, and he eventually tries to double-cross Captain Midas by attempting to take control of the derelict ship. Captain Midas discovers Spinelli’s plan and confronts him, leading to a violent altercation in which Captain Midas kills Spinelli. Throughout the story, it is clear that Captain Midas and Spinelli have a complex and troubled relationship, marked by mutual distrust and a deep-seated desire for the treasure. Their dynamic is driven by the lure of wealth and the consequences of their actions, ultimately leading to tragic consequences for both characters.
Figure 7: Examples of generated summaries in Story data using bge-large-en-v1.5 and bge-large-en-v1.5+LightPAL